OpenStudy (anonymous):

which of the following ids different from the others? a. cos^3 x-sin^3 x b. cos^4 x-sin^4 x c. 1-2sin^2 x

OpenStudy (anonymous):

\[\cos^3x -\sin^3x \]

OpenStudy (anonymous):

b. \[\cos^4x - \sin^4x\]

OpenStudy (anonymous):

c. \[1-2\sin^2x\]

OpenStudy (anonymous):

so what's the correct answer??

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

so how to find out this question. :)

OpenStudy (anonymous):

sorry, I thought I knew, but I got confused. Bump this question to the top so others can see it.